The dissolution of NH4ClO4(s) in water is endothermic, with ΔHsoln = +33.5kJ/mol.

If you prepare a 1.00 m solution of NH4ClO4 beginning with water at 21.5 ∘C, what is the final temperature of the solution (in ∘C)? Assume that the specific heats of both pure H2O and the solution are the same, 4.18 J/(K⋅g).

Express the temperature in degrees Celsius to three significant digits.
Answer #1

1.00m solution means 1 mole of NH4ClO4 is present in 1kg of water

Heat of solution= mass of water * specifc heat of water * change in temperature

33500 = 1000 * 4.184 * (change in temperature)

change in temperature = 8.0066C

Final temperature of water = 21.5 + 8.0066 = 29.5 C (in three significant digits)